OpenStudy (anonymous):

Find an equation of the tangent line to the curve at the given point x^(1/4)

OpenStudy (anonymous):

\[y(x)=x^{\frac{1}{4}}\] \[y'(x)=\frac{1}{4}x^{\frac{1}{4}-\frac{4}{4}}\] \[y'(x)=\frac{1}{4}x^{\frac{-3}{4}}\]

OpenStudy (ybarrap):

ytan = mx + b, tangent line y = x^(1/4), original m = y' = (1/4)x^(-3/4) b is determined by given point ytan = (1/4)x^(-3/4)*x + b = x^(1/4 - 3/4) + b = x^(-2/4) + b = x^(-1/2) + b Ans x^(-1/2) + b